OpenStudy (anonymous):

If f(x) = sin(x) for all x, then the average value of f on the interval [0, π] is (A) ½ (B) 1/π (C) π/2 (D) 2/π

OpenStudy (anonymous):

D

OpenStudy (anonymous):

\[1/\pi \int\limits_{0}^{\pi }\sin x dx =1/\pi (-c0sx)_{0}^{\pi }\]

OpenStudy (anonymous):

how1/pi?

OpenStudy (anonymous):

to get the average...

OpenStudy (anonymous):

why we take average??????

OpenStudy (anonymous):

because your question says...to find average...

OpenStudy (anonymous):

alright then why the average of pi????????

OpenStudy (anonymous):

the interval is given [0, pi ]

OpenStudy (anonymous):

hello??????????